The following are advertised sale prices of color televisions at an electronics store.

Size (inches) Sale Price ($)
9 137
20 187
27 297
31 447
35 1177
40 2177
60 2497

a) Calculate the least squares line. Put the equation in the form of: ŷ = a + bx.

(Round your answers to three decimal places.)

ŷ =___________+__________x

b) Find the correlation coefficient r. (Round your answer to four decimal places.)

r=____________

c) Find the estimated sale price for a 33 inch television. (Use your equation from part (d). Round your answer to two decimal places.)

$_____________

Find the estimated sale price for a 50 inch television. (Use your equation from part (d). Round your answer to two decimal places.)

$____________

d) What is the slope of the least squares (best-fit) line? (Round your answer to three decimal places.) ________________
Answer #1

From the given information, x? 20 Size Inches (x) Sale Price(y) 137 187 297 447 35 1177 40 2177 60 2497 Σx = 222 Ι Σy = 6919

α- Σy -ΣΑ 6919-54.977 (222) =-755.128949301983 =-755.129 (Round to 3 decimal place) Hence, the required least square regressi

Calculate the estimated sale price for a 33inch television. û = -755.129+54.977 (x) =-755.129+ 54.977 (33) = -755.129+1814.24

Calculate the slope of the least squares (best-fit) line. The least square regression line is: ġ=-755.129+ 54.977(x) The inte